OpenStudy (anonymous):

a^3-64a^2-4a+256 factor

OpenStudy (anonymous):

(a^2-4)*(a-64)

OpenStudy (anonymous):

=(a^2(a-64))-4(a-64) =(a^2-4)(a-64)

OpenStudy (anonymous):

=(a+2)(a-2)(a-64)

OpenStudy (anonymous):

a^2(a-64)-4(a-64) =(a-64)(a^2-4) =(a-64)(a-2)(a+2)